# Attention Guidance Mechanism for Handwritten Mathematical Expression Recognition

### 04 Mar 2024 v1

@@ 4.3 Ablation study
(1) When self-guidance is used to refine the attention maps, the ExpRates are improved on all three datasets, while the StruRates are improved slightly. This suggests that self-guidance effectively improves attention quality, but its improvement in aligning complex structures is limited due to the lack of external information.
@@ 4.4 Comparison with state-of-the-art methods
4.5 Case study with maps
This section shows some qualitative recognition results of the proposed attention guidance method. Fig. 7 illustrates the attention maps before and after attention guidance. As shown in subfig (a), the attention module without guidance fails to distinguish between the two superscripts of “$j^{2}$” and “$q^{2}$”. The attention incorrectly focuses on the image areas around “$q^{2}$” after step $t=30$. As shown in subfig (c), this is caused by the context leakage phenomenon where the middle layer mistakenly activates the areas around “$q^{2}$”. As a result, information from outside the intended context is incorporated in the attention result, thus confusing the generation process after step $t=30$. This issue is solved with the help of self-guidance and neighbor-guidance, as shown in subfig (b). Compared with subfig (a), erroneous activations on “$q^{2}$” are suppressed at step $t=31$ and $t=32$. The model can make the correct prediction based on the input context even if the erroneous activations are not entirely removed at step $t=31$. In subfig (b), the attention map at step $t$ is also the neighbor-guidance map for step $t+1$. Subfig (d) shows the guidance maps derived from the self-guidance module. As we can see, the image regions of incorrect activations in the original attention maps are assigned high weights. Using these guidance maps, erroneous correlations can be removed through the adjustment operation in Eq. 15. In conclusion, the visualization experiment shows that our proposed attention guidance method can effectively alleviate the context leakage phenomenon.
[…]
@@ 5 Conclusion
Misalignment, especially under-parsing, remains a challenging problem in HMER. We observe that one reason is the context leakage phenomenon. To address this issue, we propose an attention guidance mechanism that aims to suppress attention weights in irrelevant regions explicitly. In this way, information after the intended context is suppressed in the attention result. This paper has four main contributions: (1) We propose the context leakage phenomenon and a general attention guidance mechanism to solve this issue. (2) We devise self-guidance to refine the raw attention by seeking consensus among different attention heads since the image areas of symbols attended by multiple heads should be consistent. (3) Depending on the observation of the alignment process in the stacked decoder, we devise neighbor-guidance to reuse attention weights from the previous decoding step to refine the raw attention in the current step. (4) Experiments on standard benchmarks show that our method outperforms state-of-the-art methods, achieving ExpRates of 60.75% / 61.81% / 63.30% on the CROHME 2014 / 2016 / 2019 datasets.
The proposed attention guidance mechanism not only works for the HMER task. The ideas of self-guidance and neighbor-guidance can help improve grounding quality and alignment results for other tasks requiring dynamic alignment. In future work, we intend to apply our attention guidance mechanism to solve tasks such as document understanding, image captioning, and machine translation. Besides, we also intend to explore other attention guidance approaches to further facilitate information propagation in attention-based models.

### 05 Mar 2024 v2

@@ 4.3 Ablation study
(1) When self-guidance is used to refine the attention maps, the ExpRates are improved on all three datasets, while the StruRates are improved slightly. This suggests that self-guidance effectively improves attention quality, but its improvement in aligning expression structures is limited due to the lack of external guidance.
@@ 4.4 Comparison with state-of-the-art methods
4.5 Case study with maps
This section shows some qualitative recognition results of the proposed attention guidance method. Fig. 7 illustrates the attention maps before and after attention guidance. As shown in subfig (a), the attention module without guidance fails to distinguish between the two superscripts of “$j^{2}$” and “$q^{2}$”. The attention incorrectly focuses on the image areas around “$q^{2}$” after step $t=30$. As shown in subfig (c), this is caused by the context leakage phenomenon where the middle layer mistakenly activates the areas around “$q^{2}$”. As a result, the information behind the intended context is incorporated in the attention result, thus confusing the generation process after step $t=30$. This issue is solved with the help of self-guidance and neighbor-guidance, as shown in subfig (b). Compared with subfig (a), erroneous activations on “$q^{2}$” are suppressed at step $t=31$ and $t=32$. The model can make the correct prediction based on the input context even if the erroneous activations are not entirely removed at step $t=31$. In subfig (b), the attention map at step $t$ is also the neighbor-guidance map for step $t+1$. Subfig (d) shows the guidance maps derived from the self-guidance module. As we can see, the image regions of incorrect activations in the original attention maps are assigned high weights. Using these guidance maps, erroneous correlations can be removed through the adjustment operation in Eq. 15. In conclusion, the visualization experiment shows that our proposed attention guidance method can effectively alleviate the context leakage phenomenon.
[…]
@@ 5 Conclusion
Misalignment, especially under-parsing, remains a challenging problem in HMER. We observe that one reason is the context leakage phenomenon. To address this issue, we propose an attention guidance mechanism that aims to suppress erroneous attention explicitly. In this way, information after the intended context is suppressed in the attention result. This paper has four main contributions: (1) We propose the context leakage phenomenon and a general attention guidance mechanism to solve this issue. (2) We devise self-guidance to refine the raw attention by seeking consensus among multiple attention heads since the image regions of symbols attended by multiple heads should be consistent. (3) Depending on the observation of the alignment process of the stacked decoder, we devise neighbor-guidance to reuse attention weights from the previous decoding step to refine the raw attention in the current step. (4) Experiments on standard benchmarks show that our method outperforms state-of-the-art methods, achieving ExpRates of 60.75% / 61.81% / 63.30% on the CROHME 2014/ 2016/ 2019 datasets.
The proposed attention guidance mechanism not only works for the HMER task. The ideas of self-guidance and neighbor-guidance can help improve grounding quality and alignment results for other tasks requiring dynamic alignment. In future work, we intend to apply our attention guidance mechanism to solve tasks such as document understanding, image captioning, and machine translation. Besides, we also intend to explore other attention guidance approaches to facilitate information propagation in attention-based models further.